40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 14:12:03 2022, Yehonathan, Update, BHD, Trying LO phase locking again Screenshot_2022-09-13_17-41-33.png
    Reply  Thu Sep 15 21:12:53 2022, Paco, Update, BHD, LO phase "dc" control lophase_oltf.pdflophase_noise_spectra.pdf
       Reply  Tue Sep 27 10:50:11 2022, Paco, Update, BHD, calibrated LO phase noise lo_phase_asd.pdf
          Reply  Wed Sep 28 16:37:26 2022, Paco, Update, BHD, calibrated LO phase noise; update lo_phase_asd.pdf
             Reply  Wed Sep 28 21:54:08 2022, Paco, Update, BHD, calibrated LO phase noise; update dcpd_phase_asd.pdflo_phase_asd.pdf
                Reply  Fri Sep 30 20:18:55 2022, Paco, Update, BHD, LO phase noise with different actuation points BHDC_asd_actuation_point.png
                   Reply  Mon Oct 3 15:19:05 2022, Paco, Update, BHD, LO phase noise and control after violin mode filters filters_c1sus_lo1_lsc.pngBHDC_asd_act.pngboosted_lo_phase_control.png
Message ID: 17142     Entry time: Thu Sep 15 21:12:53 2022     In reply to: 17138     Reply to this: 17160
Author: Paco 
Type: Update 
Category: BHD 
Subject: LO phase "dc" control 

Locked the LO phase with a MICH offset=+91. The LO is midfringe (locked using the A-B zero crossing), so it's far from being "useful" for any readout but we can at least look at residual noise spectra.

I spent some time playing with the loop gains, filters, and overall lock acquisition, and established a quick TF template at Git/40m/measurements/BHD/HPC_LO_PHASE_TF.xml

So far, it seems that actuating on the LO phase through LO2 POS requires 1.9 times more strength (with the same "A-B" dc sensing). After closing the loop by FM4, and FM5, actuating on LO2 with a filter gain of 0.4 closes the loop robustly. Then, FM3 and FM6 can be enabled and the gain stepped up to 0.5 without problem. The measured UGF (Attachment #1) here was ~ 20 Hz. It can be increased to 55 Hz but then it quickly becomes unstable. I added FM1 (boost) to the HPC_LO_PHASE bank but didn't get to try it.

The noise spectra (Attachment #2) is still uncalibrated... but has been saved under Git/40m/measurements/BHD/HPC_residual_noise_spectra.xml

Attachment 1: lophase_oltf.pdf  12 kB  Uploaded Thu Sep 15 23:40:49 2022  | Hide | Hide all
lophase_oltf.pdf
Attachment 2: lophase_noise_spectra.pdf  23 kB  Uploaded Thu Sep 15 23:48:35 2022  | Hide | Hide all
lophase_noise_spectra.pdf
ELOG V3.1.3-